OpenStudy (anonymous):

whitch of the following is the slope between the points (0,-7) and (4,12)? 19/4 -19/4 4/19 -4/19

OpenStudy (mathstudent55):

The slope is the difference in the y coordinates divided by the difference in the x coordinates. The y-coordinates are 12 and -7. What is the difference between them? What is 12 - (-7) = ?

OpenStudy (anonymous):

19/4

OpenStudy (anonymous):

alright i get it thank you im not good at math at all.

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (0, -7) x1 = 0 and y1 = -7 (4,12) x 2 = 4 and y2 = 12 m = (12 - (-7) / (4 - 0) m = (12 + 7) / 4 m = 19/4 thats right

OpenStudy (texaschic101):

so does it really matter which way the formula goes ? He has the formula (y1 - y2) / (x1 - x2) I do it this way : (y2 - y1) / (x2 - x1)

OpenStudy (texaschic101):

@mathstudent55

OpenStudy (ybarrap):

It doesn't matter because \( \dfrac{-m}{-1}= m\). Or more visually: